OpenStudy (anonymous):

The radius of one sphere is twice as great as the radius of a second sphere. Find the ratio of their volumes. 1/2 1/4 1/8 1/16

OpenStudy (gorv):

volumr = 4/3*pi*r^3

OpenStudy (anonymous):

i dont get any of this at all.. :-/

OpenStudy (gorv):

4/3)*pi is constant

OpenStudy (gorv):

so volume depends on r^3 r=radius

OpenStudy (anonymous):

r1 = 3x r 2 = x a1/ a2 = 4/3)pi (r1^3)/ 4/3)pi r2^3

OpenStudy (anonymous):

so its b ? o.O

OpenStudy (gorv):

\[\frac{V1 }{ V2 }=\frac{ r1^3 }{ r2^3 }\]

OpenStudy (anonymous):

im sorry, i suck at this very bad..

OpenStudy (gorv):

lol cube 1/2^3=1/8

OpenStudy (anonymous):

lol thanks !